Operations Report – April 19th

Crew 192 Operations Report 19Apr2018

SOL 11

Name of person filing report: Richard Blakeman

Non-nominal systems: Hot water faucet in upper deck of hab, cleaned out water pump for water tank.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 61% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can.

Water (trailer) – Empty water tank secure ob Trailer.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – Yes
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133755.8

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas full.
Assumption on percentage …untenable to check
Deimos rover used: No
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A great EVA was followed by cleaning and preparing the hab for Crew 193.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine. Inventoried and cleaned 30 ear pieces not including 5 our crew has.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Thoroughly cleaned and prepared for next crew.

Summary of RAM operations: Trailer water tank moved out of RAM and back onto trailer and secured. Doctor requests water test kit before use. Cleaned, swept and mopped RAM…secured doors.

Summary of health and safety issues: Crew was very tired after working on water tank, so afternoon EVA was changed to an engineering EVA.

Questions, concerns, and requests to Mission Support: Fuel for ATVs

Copyright © The Mars Society. All rights reserved. | Main Site